Verify Euler's realtion for

z= tan(y/x) ,x ≠ 0


1
Expert's answer
2021-05-07T14:31:49-0400

We have given that,

"z = f(x,y) = tan\dfrac{y}{x}" "x \ne 0"


We have to prove that,

"x \dfrac{\delta f}{\delta x} + y \dfrac{\delta f}{\delta y} = -f"


Solving the LHS


"x \dfrac{\delta f}{\delta x} + y \dfrac{\delta f}{\delta y} = -\dfrac{y}{x}sec^2\dfrac{y}{x}+\dfrac{y}{x}sec^2\dfrac{y}{x} = 0"


which is not equal to RHS. Hence, Euler's relation is not valid.